# UniLiP: Adapting CLIP for Unified Multimodal Understanding, Generation and Editing

### v2

@@ 1 Introduction
Early efforts (Team, 2024; Zhou et al., 2024) combine VQVAE or VAE with LLMs for unification, but often achieve subpar performance in understanding because the encoder struggles to capture rich semantics. To mitigate this, recent works build unified tokenizers based on CLIP (Wu et al., 2024; Sun et al., 2024b; Qu et al., 2024; Han et al., 2025). While CLIP features possess rich semantics, they lack pixel details, requiring extra modifications for reconstruction. As depicted in Figure 1 (a), VILA-U et al. (Wu et al., 2024; Qu et al., 2024) discretize CLIP features, at the cost of information loss and worse understanding than the original CLIP. Emu2 (Sun et al., 2024b) freezes CLIP and finetunes a diffusion decoder to reconstruct images from CLIP features. However, as CLIP features lose pixel details, the diffusion decoder cannot produce consistent outputs (example in Figure 1), hindering editing tasks. Therefore, these methods either compromise CLIP’s original understanding performance or achieve inconsistent reconstruction, leading to a trade-off in unification.
@@ 2 Related Work
MLLMs for Image Understanding. Initial approaches (Liu et al., 2023; Zhu et al., 2023; Dai et al., 2023) align the features of visual encoders to LLM input spaces, achieving strong performance through instruction tuning. In this architecture, the visual encoder is crucial for perception, and CLIP (Radford et al., 2021) has become the optimal choice due to its robust alignment with text. Subsequent methods have explored the use of better CLIP, including SigLIP (Zhai et al., 2023) and InternViT (Chen et al., 2024d), enhancing performance through improved designs (Shi et al., 2024a; Tong et al., 2024a), larger model sizes (Liu et al., 2024a), better training strategies (Deitke et al., 2024), and extension to fine-grained perception tasks (Wang et al., 2024b; Tang et al., 2025). Nevertheless, these models remain specialized in comprehension without generative capabilities.
Unified MultiModal Understanding and Generation. Early efforts (Team, 2024; Zhou et al., 2024) typically share model weights and employ autoregressive modeling based on VQ-VAE or diffusion modeling utilizing VAE. However, these approaches often struggle to resolve task conflicts and suffer from shallow semantics. Subsequent works generally follow three main directions: First, decoupling the image encoder (Chen et al., 2025c) or introducing mixture-of-experts (MoE) (Shi et al., 2024b; Deng et al., 2025; Li et al., 2025; Liao et al., 2025) helps reduce task conflicts and improve model performance, but this benefit comes at the cost of increased complexity. Second, quantizing CLIP features (Wu et al., 2024; Qu et al., 2024; Huang et al., 2025; Ma et al., 2025) can boost understanding, though still lag behind with continuous CLIP-based models. Third, an increasingly popular strategy combines frozen MLLM with DiT (Tong et al., 2024b; Pan et al., 2025; Lin et al., 2025a; Wu et al., 2025b). Recently, BLIP3-o (Chen et al., 2025a) utilizes a DiT to generate CLIP features, achieving better prompt alignment than VAE-based approaches. Nonetheless, BLIP3-o faces two issues: (1) Following the reconstruction paradigm of Emu2 (Sun et al., 2024b), it fails to guarantee consistent reconstruction and editing. (2) Image generation requires two diffusion processes, first for CLIP features and then for VAE features. In contrast, UniLIP explicitly supplements the image details missing in CLIP features, enabling excellent reconstruction, generation, and editing performance through a lightweight pixel decoder and a single diffusion process.
@@
Instruction-Guided Image Editing. Early methods (Brooks et al., 2023; Zhang et al., 2023) employ modular pipelines, using MLLMs to generate prompts or spatial cues that guide editing. To enhance semantic control, a significant trend is to explore direct latent-level fusion between MLLMs and diffusion decoders. Many approaches (Liu et al., 2025; Wu et al., 2025a) utilize MLLMs to extract rich semantic embeddings from both image and text, thereby conditioning the generative backbone for more faithful and controllable edits. However, since CLIP in MLLMs cannot capture pixel details, these methods still require a VAE for supplementary. UniWorld-V1 (Lin et al., 2025a) finds that SigLIP can preserve consistency at high resolution, but still relies on VAE features for generation. In contrast, UniLIP can achieve consistent editing by adapting CLIP for reconstruction, maintaining coherence without resolution constraints and VAE dependence.
@@ 3.1 Finetuning CLIP for Image Reconstruction
where $\lambda$ is the weight for the distillation loss, $F_{\text{orig}}$ denotes the original CLIP features, and $F_{\text{ft}}$ is the fine-tuned CLIP features. We empirically find that setting $\lambda$ to 1 is sufficient (see Table 11).
@@ 3.2 Dual conditions for image generation and editing
As UniLIP produces continuous features, we follow methods like MetaQuery (Pan et al., 2025) and BLIP3-o (Chen et al., 2025a) to build the generation pipeline. These methods employ query embeddings to bridge a frozen MLLM with a diffusion transformer, where the query tokens serve as the conditions for the generation process. Notably, in knowledge-augmented generation, query embeddings encode the MLLM’s reasoning result to implicitly define the target image content, resulting in better prompt alignment. However, such methods are rarely applied to editing tasks and have not demonstrated satisfactory quality or robust quantitative metrics in editing benchmarks.
We identify the main bottleneck as the fixed number of query embeddings. For example, BLIP3-o uses 64 tokens and MetaQuery uses 256 tokens, which limit the expressive capacity of the queries. This problem is not obvious in text-to-image generation, since most text prompts in generation benchmarks are short and LLMs excel at compressing text information. However, in editing, query embeddings need to retain the details of one or more reference images. Therefore, using a fixed number of tokens is inevitable for information loss, which leads to inconsistency in the editing.
@@ 4.1 Implementation Details
Image Generation and Editing Training. We adopt a three-stage training approach. In the first stage, we freeze the MLLM and diffusion transformer, and train only the connector on generation data for 50k steps. In the second stage, we train both the connector and diffusion transformer on generation and editing data for 200k steps. In the third stage, we train on instruction-tuning data for both generation and editing tasks for 30k steps. All stages use a batch size of 512 with a cosine learning rate decay schedule ranging from 1e-4 to 1e-5.
@@ B.3 Inference Speed
Table 11: Ablation of distillation loss types and weights.
@@ C.3 Distillation Loss
Table 11 ablates the distillation loss type and weight. At a fixed weight, MSE is most effective for preserving understanding, while cosine loss is least effective because it only penalizes differences in feature direction. Regarding weighting, low values are detrimental to understanding, yet high values offer no significant gains, likely due to overfitting that degrades generalization and reconstruction. Therefore, MSE with a weight of 1.0 achieves the optimal trade-off.

### v3

@@ 1 Introduction
Early efforts (Team, 2024; Zhou et al., 2024) combine VQVAE or VAE with LLMs for unification, but often achieve subpar performance in understanding because the encoder struggles to capture rich semantics. To mitigate this, recent works build unified tokenizers based on CLIP (Wu et al., 2024b; Sun et al., 2024b; Qu et al., 2024; Han et al., 2025). While CLIP features possess rich semantics, they lack pixel details, requiring extra modifications for reconstruction. As depicted in Figure 1 (a), VILA-U et al. (Wu et al., 2024b; Qu et al., 2024) discretize CLIP features, at the cost of information loss and worse understanding than the original CLIP. Emu2 (Sun et al., 2024b) freezes CLIP and finetunes a diffusion decoder to reconstruct images from CLIP features. However, as CLIP features lose pixel details, the diffusion decoder cannot produce consistent outputs (example in Figure 1), hindering editing tasks. Therefore, these methods either compromise CLIP’s original understanding performance or achieve inconsistent reconstruction, leading to a trade-off in unification.
@@ 2 Related Work
MLLMs for Image Understanding. Initial approaches (Liu et al., 2023; Zhu et al., 2023; Dai et al., 2023) align the features of visual encoders to LLM input spaces, achieving strong performance through instruction tuning. In this architecture, the visual encoder is crucial for perception, and CLIP (Radford et al., 2021) has become the optimal choice due to its robust alignment with text. Subsequent methods have explored the use of better CLIP, including SigLIP (Zhai et al., 2023) and InternViT (Chen et al., 2024d), enhancing performance through improved designs (Shi et al., 2024; Tong et al., 2024a), larger model sizes (Liu et al., 2024a), better training strategies (Deitke et al., 2024), and extension to fine-grained perception tasks (Wang et al., 2024b; Tang et al., 2025). Nevertheless, these models remain specialized in comprehension without generative capabilities.
Unified Visual Tokenizers. Early works use VQ/VAE as unfied encoders (Zhou et al., 2024), which suffer from poor understanding due to limited semantics. Subsequent methods decouple the encoder, using CLIP for understanding and VQ/VAE for generation (Wu et al., 2024a), which improves performance but increases complexity. To simplify this, recent approaches build unified encoders upon CLIP. Some quantize CLIP features for reconstruction (Wu et al., 2024b; Qu et al., 2024; Ma et al., 2025a), causing semantic degradation. Other methods freeze CLIP and train a diffusion decoder to generate images conditioned on CLIP features (Sun et al., 2024b). However, since CLIP features lose image details, the reconstruction quality is poor. In contrast, UniLIP can maintain CLIP’s understanding capability while achieving excellent reconstruction.
Architectures for Unified MultiModal Understanding and Generation. Early methods share weights between tasks (Team, 2024), leading to training instability. Later methods adopt Mixture-of-Experts (MoE) to decouple tasks, resolving conflicts but doubling parameters (Liao et al., 2025; Deng et al., 2025). To improve integration, the pioneering work DreamLLM (Dong et al., 2023) first proposes learnable queries to bridge MLLMs and diffusion models, facilitating synergistic joint training. MetaQuery (Pan et al., 2025) further extends this to connect frozen MLLMs with DiTs. BLIP3-o (Chen et al., 2025a) adopts a similar framework but uses CLIP features as diffusion targets for better prompt alignment. While UniLIP also uses learnable queries, we introduce a dual-condition design to address the inability of prior query-based methods in editing.
@@
Instruction-Guided Image Editing. Early methods (Brooks et al., 2023; Zhang et al., 2023) employ modular pipelines, using MLLMs to generate prompts or spatial cues that guide editing. To enhance semantic control, many approaches (Liu et al., 2025; Wu et al., 2025a) utilize MLLMs to extract rich semantic embeddings from both image and text, thereby conditioning the generative backbone for more faithful and controllable edits. However, since CLIP in MLLMs cannot capture pixel details, these methods still require a VAE for supplementary. UniWorld-V1 (Lin et al., 2025a) finds that SigLIP can preserve consistency at high resolution, but still relies on VAE features for generation. In contrast, UniLIP can achieve consistent editing by adapting CLIP for reconstruction, maintaining coherence without resolution constraints and VAE dependence.
@@ 3.1 Finetuning CLIP for Image Reconstruction
where $\lambda$ is the weight for the distillation loss, $F_{\text{orig}}$ denotes the original CLIP features, and $F_{\text{ft}}$ is the fine-tuned CLIP features. We empirically find that setting $\lambda$ to 1 is sufficient (see Table 11). Additionally, to further constrain parameter updates in CLIP, we set its learning rate to 0.1 times the global learning rate. We find that this strategy also helps preserve original abilities (Table 6).
@@ 3.2 Dual conditions for image generation and editing
As UniLIP produces continuous features, we follow methods like DreamLLM (Dong et al., 2023) and MetaQuery (Pan et al., 2025) to build the generation pipeline. These methods employ query embeddings to bridge MLLM with a diffusion transformer, where the query tokens serve as the conditions for the generation process. Notably, in knowledge-augmented generation, query embeddings encode the MLLM’s reasoning result to implicitly define the target image content, resulting in better prompt alignment. However, such methods are rarely applied to editing tasks and have not demonstrated satisfactory quality or robust quantitative metrics in editing benchmarks.
We identify the main bottleneck as the fixed number of query embeddings. For example, DreamLLM uses 64 tokens and MetaQuery uses 256 tokens, which limit the expressive capacity of the queries. This problem is not obvious in text-to-image generation, since most text prompts in generation benchmarks are short and LLMs excel at compressing text information. However, in editing, query embeddings need to retain the details of one or more reference images. Therefore, using a fixed number of tokens is inevitable for information loss, which leads to inconsistency in the editing.
@@
3.3 Training Schemes
To construct a unified model based on UniLIP capable of simultaneous understanding, generation, and editing, we adopt a three-stage training strategy:
Stage 1: Connector Training. The primary goal of this stage is to train the connector to effectively bridge the MLLM and the DiT. The connector learns to align the MLLM’s output features with the DiT’s original conditional feature space. Parameters of both the MLLM and the DiT are frozen in this stage, and training is conducted exclusively on generation tasks.
Stage 2: Unified Generation and Editing Training. We utilize large-scale datasets to train the model on general generation and editing. Both the connector and the DiT are trained in this stage. Consistent with prior methods (Chen et al., 2025a), the MLLM remains frozen, eliminating the need for expensive training on understanding tasks.
Stage 3: Supervised Fine-Tuning (SFT). We employ high-quality instruction tuning data to enhance generation fidelity and prompt alignment for both generation and editing tasks. During this phase, the connector and DiT continue to be optimized, while the MLLM remains frozen.
@@ 4.1 Implementation Details
Image Generation and Editing Training. Stages 1, 2, and 3 are trained for 50k, 200k, and 30k steps, respectively. All stages use a batch size of 512 with a cosine learning rate decay schedule ranging from 1e-4 to 1e-5.
@@ 4.3 Ablation Study
Two stage vs. Single stage. To verify the effectiveness of the two-stage training strategy, we compare the training loss and performance of our approach against a single-stage baseline in Figure 6. We observe that the single-stage method exhibits severe instability, characterized by an initial distillation loss spike nearly double that of our method (0.0939 vs. 0.0497). Consequently, its convergence is significantly slower, requiring $3\times$ more iterations to reduce the loss and $4\times$ more to recover understanding performance. These results suggest that Stage 1 is critical for pre-aligning the pixel decoder; without this preliminary step, the initial misalignment between the unfrozen CLIP and the randomly initialized projection layer leads to gradient instability.
Figure 6: Comparison of training loss and performance curve between single-stage and two-stage method. Thanks to the pre-alignment in Stage 1, our Stage 2 can converge much faster.
@@ 5 Conclusion
Acknowledgments
LW is supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005).
@@ B.3 Inference Speed
B.4 Reasoning Editing
To comprehensively assess UniLIP’s editing capabilities, we evaluated it on the RISE benchmark. Results in Table 13 show that our 3B model outperforms significantly larger models like Step1X-Edit (7B+12B) and Bagel (7B+7B), confirming UniLIP’s superiority in reasoning-based editing.
B.5 Linear and attentive probe
We conduct both linear probe and attentive probe evaluations on ImageNet. The results in Table 13 show that our UniLIP achieves classification performance comparable to the original InternViT. This validates the effectiveness of our reconstruction training and self-distillation loss.
@@ B.5 Linear and attentive probe
Table 11: Ablation of distillation loss types and weights. “Cls” means classification loss.
Model Size Temporal Causal Spatial Logic Overall
OmniGen (Xiao et al., 2025) 3.8B 1.2 1.0 0.0 1.2 0.8
Step1X-Edit (Liu et al., 2025) 7B + 12B 0.0 2.2 2.0 3.5 1.9
BAGEL (Deng et al., 2025) 7B + 7B 2.4 5.6 14.0 1.2 6.1
[…]
Table 12: Evaluation of image editing ability on RISE benchmark.
Model Linear Attentive
Frozen CLIP 79.8 82.6
[…]
Table 13: Linear and attentive probing results on ImageNet validation.
@@ C.3 Distillation Loss
Table 11 ablates the distillation loss type and weight. At a fixed weight, MSE is most effective for preserving understanding, while classification loss is least effective because it provides only coarse, image-level supervision signals. Regarding weighting, low values are detrimental to understanding, yet high values offer no significant gains, likely due to overfitting that degrades generalization and reconstruction. Therefore, MSE with a weight of 1.0 achieves the optimal trade-off.
C.4 Understanding the Query Embeddings
We employ learnable queries primarily to enable knowledge-augmented generation (e.g., for WISE benchmarks). In such tasks, prompts like “Draw Thanksgiving food” do not explicitly describe the visual details, so the model must rely on internal knowledge to deduce what to draw. To better illustrate the effect of query embeddings, we treat the output query embeddings as text embeddings and decode them into text. We observe that the decoded text can explicitly identify the target content. For instance, given the prompt “Thanksgiving food”, the decoded text includes the word “turkey”, which greatly simplifies generation for the DiT. In contrast, models lacking this mechanism generate incorrect foods. We provide detailed visual comparisons in Figure 8.
Figure 8: Qualitative comparison between multimodal hidden states and query embeddings. We decode query embeddings back into text (masking unreadable tokens with …). The results show high semantic consistency with both the prompt and the generated image, especially the red-highlighted terms.
C.5 Query length
We investigate the impact of the learnable query length, denoted as $L$, within our dual-condition architecture. As presented in Table 14, the performance on the WISE benchmark exhibits a positive correlation with query length, reaching a point of convergence around $L=256$. We hypothesize that longer queries effectively mimic an increased inference budget for textual reasoning. This expanded capacity enables the Large Language Model (LLM) to more comprehensively activate its internal knowledge, thereby facilitating more accurate image generation.
Query Length Cultural Time Space Biology Physics Chemistry. Overall
0 0.43 0.50 0.62 0.44 0.58 0.35 0.47
64 0.48 0.52 0.62 0.42 0.54 0.40 0.50
128 0.52 0.56 0.68 0.48 0.58 0.45 0.54
256 0.54 0.58 0.70 0.50 0.62 0.46 0.56
512 0.56 0.57 0.67 0.48 0.60 0.44 0.56
Table 14: Ablation study on the impact of learnable query length on the WISE benchmark.
C.6 Joint Training
We unify image generation and editing to exploit task synergy, specifically enabling generation capabilities to facilitate editing tasks. Given the scarcity of high-quality editing data (1.5M samples) compared to generation data (36M), training solely on the former poses a significant risk of overfitting. By incorporating large-scale generation data, the model captures a more robust underlying image distribution, leading to more realistic and consistent edits. To validate this, we train a 1B-parameter variant using only the editing dataset. As demonstrated in Table 15, joint training significantly outperforms the “Edit Only” baseline across all metrics, confirming that generative knowledge is crucial for high-quality editing.
Method Add Adj. Ext. Repl. Rmv. Bkg. Style Hyb. Act. Overall
Edit Only 3.99 3.38 1.77 4.16 3.92 3.88 4.75 2.96 4.21 3.67
Joint Training 4.11 3.58 1.94 4.30 3.97 4.00 4.87 3.17 4.35 3.81
Table 15: Ablation study on training strategies. We compare a model trained solely on editing data against our joint training strategy. The joint approach leverages large-scale generation data to enhance editing performance across various sub-tasks.
C.7 Comparison with Dual-Encoder methods
To demonstrate the superiority of our unified encoder over the dual-encoder architecture, we conduct ablation studies based on JanusFlow (Ma et al., 2025b). We construct a unified variant, UniLIP-JanusFlow, by replacing the original dual encoders (SigLIP and SDXL-VAE) with our UniLIP-SigLIP. Specifically, we first apply our two-stage reconstruction training to the SigLIP encoder within JanusFlow to enable reconstruction (UniLIP-SigLIP). Subsequently, using JanusFlow’s pre-trained weights, we fine-tune both the original dual-encoder baseline (JanusFlow) and our unified model (UniLIP-JanusFlow) on the ShareGPT-4o-Image Chen et al. (2025b) dataset for 10k steps.
We first evaluate the fundamental capabilities of the encoders as shown in Table 16. In terms of reconstruction, UniLIP-SigLIP achieves a lower rFID (0.35) than the original SDXL-VAE (0.38) while maintaining comparable PSNR and SSIM, proving it can serve as a high-quality visual generator. Crucially, this added capability does not compromise semantic understanding; UniLIP-SigLIP preserves or slightly improves performance on multimodal benchmarks (e.g., MME-P, MMVP) compared to the original SigLIP. This confirms that our approach effectively resolves the conflict between reconstruction and understanding tasks.
Reconstruction Understanding
[…]
JanusFlow (Ma et al., 2025b) 0.38 26.79 0.827 1302.0 64.6 65.0 65.7 55.5
[…]
Table 16: Reconstruction and understanding performance comparisons.
Next, we compare the performance on generation and editing tasks (Table 17). To extend JanusFlow for the editing task, we follow previous work (Chen et al., 2025b) by concatenating SigLIP and VAE features, resulting in a total of 1152 image tokens. In contrast, UniLIP-JanusFlow utilizes a single unified feature sequence of only 576 tokens. Despite the significant reduction in token count, UniLIP-JanusFlow consistently outperforms the dual-encoder baseline (+2.0 on GenEval, +5.0 on WISE, and +0.14 on ImgEdit). These results demonstrate that the unified representation eliminates the feature misalignment inherent in dual-encoder systems and provides a more efficient and effective foundation for generative tasks.
Model GenEval WISE ImgEdit
Sgl. Obj. Two Obj. Overall Time Space Overall Add Repl. Overall
JanusFlow (Ma et al., 2025b) 0.99 0.82 0.73 0.35 0.46 0.31 3.21 2.92 2.95
[…]
Table 17: Generation and Editing Performance. Models are fine-tuned on ShareGPT-4o-Image. We report representative sub-metrics (e.g., Single/Two Object for GenEval, Time/Space for WISE, Add/Replace for ImgEdit) alongside the overall scores. UniLIP-JanusFlow consistently outperforms the dual-encoder baseline across both fine-grained and aggregate metrics.